Check my workCheck My Work button is now enabled

Item 5

Item 5

Observations are taken on sales of a certain mountain bike in 24 sporting goods stores. The regression model was Y = total sales (thousands of dollars), X1 = display floor space (square meters), X2 = competitors' advertising expenditures (thousands of dollars), X3 = advertised price (dollars per unit).

(a)

Fill in the values in the table given here. (Negative values should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Round your t-values to 3 decimal places and p-values to 4 decimal places.)

   

  Predictor Coefficient SE      tcalc      p-value
  Intercept 1,265.1     343.6           
  FloorSpace 11.455     1.54           
  Competing Ads -6.453     3.908           
  Price -0.14956     0.08880           
Answer #1

t value = Coefficient / SE

eg: t value for Price = -0.14956/0.0888 = -1.684

degrees of freedom (df) = n-2 = 24-2 = 22

p-value = TDIST(abs(t value),df,2)

eg: p-value of Price = TDIST(1.684,22,2) = 0.1063

Hence, Completed Table:

  Predictor

Coefficient

SE

     tcalc

     p-value

  Intercept

1,265.10

343.6

3.682

0.0013

  FloorSpace

11.455

1.54

7.438

0.0000

  Competing Ads

-6.453

3.908

-1.651

0.1129

  Price

-0.14956

0.0888

-1.684

0.1063

PS: If Satisfied, kindly upvote